experiment result:

2 drops of approximately 0.1M NaOH + BTB were titrated by 25 drops of 0.10M HCl.

Question:

Show your calculations for the determination of the concentration of the NaOH solution. Report your value to 3significant figures
Answer #1

The concentration of the NaOH solution is already given in question is 0.10M. I think it entered by mistake.

But ignoring the given concentration of NaOH solution 0.10M, we can calculate the same from other results.

Volume of NaOH solution (V1) = 2 drops

Volume of HCl(V2) = 25 drops

Concentration of HCl (M2) = 0.10M

Concentration of NaOH solution(M1) = ?

We know the that \small M1V1=M2V2

\small M1=\frac{M2V2}{V1}

\small M1=\frac{25*0.10}{2} = 1.25M

Therefore the concentration of NaOH solution = 1.25M